OpenStudy (anonymous):

pls help me to find out function's domain and range. Next, describe level curves of the functions i) f(x,y)= x-y ii) f(x,y) = x^2+y^2 iii) f(x,y) = x^2-y^2 in square root iv) f(x,y,z) = x+y+z

OpenStudy (anonymous):

So you need to state the domain and range, as well as a description of the level curves of each function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Have you gone through any of these yet? If not, we can start with the first one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok so let's do the first one then?

OpenStudy (anonymous):

So when determining the domain of any given function, we need to describe the set of all values of the independent variables. For i ), would you say that there are any restrictions on the domain?

OpenStudy (anonymous):

i think no

OpenStudy (anonymous):

Mmmmmk that's correct. And what's your reasoning behind that?

OpenStudy (anonymous):

zero denominators and negatives under square roots

OpenStudy (anonymous):

Let's consider the range now.

OpenStudy (anonymous):

Wait. . .

OpenStudy (anonymous):

When a real-valued function is defined along all values of the x-axis, we say that its domain is R, or, the set of all real numbers.

OpenStudy (anonymous):

Likewise, we can ask ourselves if there are any restrictions upon the range.

OpenStudy (anonymous):

In other words, what is the set of all values of f(x,y)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then, like the domain, the range is all real numbers.

OpenStudy (anonymous):

0 to infinity

OpenStudy (anonymous):

I would just write something like dom f(x,y) = R and ran f(x,y) = R

OpenStudy (anonymous):

Or, in interval notation, from (-inf, inf)

OpenStudy (anonymous):

Do you know what notation I am referring to when I write R?

OpenStudy (anonymous):

real number

OpenStudy (anonymous):

Yes, all real numbers. OK. What do you think the dom and ran of ii ) is?

OpenStudy (anonymous):

same

OpenStudy (anonymous):

Mmmmmm, not quite.

OpenStudy (anonymous):

The domain is the same. Both x and y can be any real number. However, what's the lower bound of the range? Is the range ever negative?

OpenStudy (anonymous):

no..from 0 to infinity

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Oh, we forgot the level curves. Ok, real quick, then I have to go. For i ), if we let x - y = c , where c is an arbitrary constant, then we derive a level curve of the surface in three-space. Start by letting c = 0.

OpenStudy (anonymous):

What kind of equation do you derive from doing this and along what value of f(x,y) does the image lie on?

OpenStudy (anonymous):

functions of several variables and the graph?

OpenStudy (anonymous):

Here, if we let x- y = 0, then it follows that y = x and so we have a line that lies along the plane f(x, y) = 0.

OpenStudy (zzr0ck3r):

I think technically the domain is \[R\times R\] and the range is \(R\)

OpenStudy (zzr0ck3r):

for the first one anyway...

OpenStudy (anonymous):

I think you are right. Since there are two independent variables.

OpenStudy (zzr0ck3r):

well functions take an input, and two different elements would not be 1 input. So we are are looking at f((x,y)) where the input is an element of some set of ordered pairs.

OpenStudy (zzr0ck3r):

Anyway sorry to interrupt its a small distinction, and one that maybe the teacher would not care about:)

OpenStudy (anonymous):

Yes, thank you for clarifying. I'm like thinking half between R and R X R. . .

OpenStudy (anonymous):

I must go. Good night and good luck. I think you have the general idea as far as the dom and ran go. You can go to Pauls Online Notes or the Khan Academy for more info regarding this topic.

OpenStudy (anonymous):

ok..tq for your kindness reply

OpenStudy (anonymous):

pls help me to solve this Use the method of Lagrange multipliers to find the local extreme values of the function f subject to the constraint: We introduce the multiplier λ and solve the system ∇=∇ f g λ , g = 0. i. f x y xy (, ) = on the ellipse 2 2 x y + = 4 1 ii. f xy x y (, ) 2 6 = −+ on the circle 2 2 x y + = 4

OpenStudy (anonymous):

Use the method of Lagrange multipliers to find the local extreme values of the function f subject to the constraint: We introduce the multiplier λ and solve the system ∇f=λdelta g , g = 0. i. f( x, y)= xy on the ellipse x^2+4y^2= 1 ii. f (x,y)=2x- y+ 6 on the circle x^2 + y^2= 4

OpenStudy (anonymous):

okay, for (i) we know \(\nabla f=(y,x)\) whereas \(\nabla g=(2x,8y)\) so:$$y=2\lambda x,x=8\lambda y$$and thus \(y=2\lambda(8\lambda y)=16\lambda ^2y\) meaning either \((16\lambda^2-1)y=0\) so \(x=y=0\) or \(\lambda=\pm1/4\). since \(x=y=0\) does not satisfy the constraint we look substitute in for \(\lambda\) to get \(x=\pm2y\):$$x^2+4y^2=1\\(\pm2y)^2+4y^2=1\\4y^2+4y^2=1\\8y^2=1\\y=\pm\frac1{2\sqrt2}$$thus \(x=\pm1/\sqrt2\) and we get four extrema